You are here: Foswiki>Commissioning Web>Com20090629 (30 Jun 2009, TonyEdgin)Edit Attach

UT20090629 Telescope Performance Verification
with LUCIFER at LFBG

Observers: W. Seifert, N. Ageorges, V. Knierim, J. Hill (LBT)
Telescope Operator: A. Ceranski
Software Support: T. Edgin (Tucson)
Instrument Support: T. Shih
Telescope Support: B. Curtis

Summary

Most of night lost to clouds and the threat of rain. Opened at 02:50 AM (MST) to clear sky, but then the clouds moved back in on us. Used 1 hour to collect data to check pointing.

Preparations

Details

02:40 Open vent doors, and run mirror ventilation. Sky is full of clouds of varying thickness. Storms to the south are dying away.

05:40 Close vent doors

07:00 Small thunderstorm approaching from NW.

Open

09:50 Open with T=11.1 degC, D=6.7 degC, sky 80% clear (really clear), with wind 4 m/sec from W.

10:00 preset TRACK PARALLACTIC

10:03 Restart PCS to get polys flowing for first preset

10:04 preset again - Adjust IE=-96.8 CA=-41.68

10:11 preset ACTIVE - seeing is 0.7-0.8 arcsec after a few active iterations

10:19 preset TRACK again to verify pointing as we didn't recognize previous field. Pointing is OK, but sky has a uniform layer of cloud again. Adjust CA=-31.68.

Pointing check through the clouds

Note that telescope is not really in thermal equilibrium, so this data is more of a check than the basis of a new pointing model. C-Rings are about 0.8 degC too warm.

10:23 turn on pointing log = ptlog090629

10:25 Save image of the previous star - luci

10:27 preset TRACK POSITION - to the same star ACT0272 - luci090628 003

10:30 preset TRACK 22:19:07 +28:50:51 WT10_306 - luci 005

10:32 preset TRACK 21:54:35.823 -02:59:49.5 WT10_202 - luci 007 (after waiting for cloud to pass)

10:37 preset TRACK ACT0265 - luci 009

10:40 preset TRACK 0070 - too low for these conditions, plus primary hits the Z/RX limit

10:45 preset TRACK WT10_115 - luci 010

10:46 preset TRACK ACT0166 - luci 011

10:50 preset TRACK ACT0238 - luci 012 (saturated)

10:53 preset ACTIVE - didn't find a star. Adjust pointing IE=-104.3 CA=-54.43

10:59 preset ACTIVE - variable clouds cause us to lose guide star briefly. GCS manages the WFS images OK in this sense - it knew when to not do WFS corrections.

11:04 preset TRACK WT10_349 - luci 013 (saturated)

11:08 preset TRACK

11:10 preset TRACK ACT0272 - luci 014

11:12 preset TRACK ACT0219 - luci 015

11:13 preset TRACK ACT0183 - luci 016

11:15 preset TRACK ACT0382 - luci 017 (crowded image)

Close

11:18 Close for new radar echo and twilight. T=11.9 degC D=3.2 degC Wind 6 m/sec fron NW.

11:19 stop pointing log

The foam on the web camera microphone does help with the wind noise. We may need another layer in high wind.

Note that SMT ambient reads 10.4 degC, but the ambient temperature in telemetry goes below +10.0 degC for the last hour. It seems like the number in telemetry does not match what is displayed on the weather GUI. ???

AFAIK, the SMT weather data are not being written to telemetry. --Tony

-- JohnHill - 29 Jun 2009
Topic revision: r4 - 30 Jun 2009, TonyEdgin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback